OpenStudy (anonymous):

Change -5x + 2y = 12 into slope-intercept form. Choose one answer. a. y = 5x + 6 b. y = 5/2x+6 c. y = -5/2x+6 d. y = 5x - 6

OpenStudy (anonymous):

answer b

OpenStudy (angela210793):

2y=12+5x y=12/2+5x/2 y=6+5x/2 B

OpenStudy (anonymous):

2y=5x+12 y=5/2x+6

OpenStudy (anonymous):

thanks! thats what i was thinking. thanks for the help :)

OpenStudy (anonymous):

ma pleasure!!!!!!